# Yoga for Fourier–Mukai partnership

### v2

@@ Abstract.
We study the behavior of integral transforms under various forms of base change. In particular, we establish a yoga of local algebra and fibers to test for derived equivalences via integral transforms. This partially generalizes a result of Orlov to singular varieties and strengthens many results in the literature to arbitrary fields. Additionally, it provides new insight into fibrations and their singularities arising in more arithmetically flavored contexts (e.g. projective and flat schemes over a DVR).
@@
Fourier–Mukai partners, integral transforms, singular varieties, derived categories
@@ 1.1. What is known
The context of our work starts with Mukai [Muk81]. Loc. cit. introduced integral transforms to study triangulated equivalences between derived categories associated with abelian varieties. Over time, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g. [Bri02, Kaw02]), with further applications in string theory [AHR05, ASS05].
Much of this progress has focused on smooth varieties, where geometric results rely on categorical methods. However, extending these methods to singular varieties remains challenging, largely because bounded coherent complexes need not be perfect. Thus, an important task is to develop techniques that improve our understanding of integral transforms in such singular settings.
Recent work has begun to address these difficulties. Criteria determining when an integral transform is fully faithful or an equivalence have been established for varieties over algebraically closed fields, generalizing results from the smooth case [HLS07, HLS09] and building upon foundational work of Bondal–Orlov [AB95]. Likewise, Bondal–Orlov’s reconstruction theorem for smooth projective varieties has been extended to projective curves [Spe23] and (anti-)Fano Gorenstein varieties [Bal11].
There is also growing interest in studying integral transforms in the relative setting. Specifically, let $Y_{1}$ and $Y_{2}$ be proper $S$-schemes where $S$ is Noetherian. An integral transform is an exact functor $\Phi_{K}:=\mathbf{R}(p_{2})_{\ast}(\mathbf{L}p_{1}^{\ast}(-)\otimes^{\mathbf{L}}K)\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ where $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$ and $p_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ are the projection morphisms. If $\Phi_{K}$ is an equivalence, then we say $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners (over $S$).
Such transforms naturally arise in the study of fibrations; that is, in situations involving proper and flat morphisms. See e.g. [HLS07, HLS09], which dealt with families whose special fibers have mild singularities. In particular, loc. cit. investigated when an integral transform between locally projective fibrations $Y_{i}\to S$ is an equivalence for $S$ of finite type over an algebraically closed field. It was shown that the induced integral transform on special fibers (i.e. closed points) over $S$ controls the behavior of equivalences and full faithfulness.
@@ 1.2. What we do
In this work, we deepen this perspective by systematically exploiting special fibers. Specifically, we study the behavior of integral transforms under base change in settings that incorporate local algebra. Moreover, our approach naturally extends several results in the literature to arbitrary fields, and at the same time yields new insight for singularities of fibrations.
@@
Neeman has developed a new framework called approximable triangulated categories (see e.g. [CNS25, Nee23]), which has resolved open conjectures in algebraic geometry [Nee21a, Nee24]. As a consequence of this machinery, any integral transform $\Phi_{K}$ inducing an equivalence on $D_{\operatorname{qc}}$ restricts to equivalences on the subcategories of bounded complexes with coherent cohomology $D^{b}_{\operatorname{coh}}$ and perfect complexes $\operatorname{Perf}$. Moreover, an equivalence on these ‘smaller’ categories implies an equivalence on the ‘larger’ category. See Remark˜2.7 for details.
@@
Consequently, questions about equivalences may be reduced to the behavior of integral transforms on these smaller subcategories. Whether an integral transform restricts to these smaller categories is now well-understood [Riz17, Bal09, DLM25]. These results shed light on the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
@@
However, the behavior of adjointness under base change is not yet fully understood. Our focus is on base changes induced by affine morphisms of finite tor-dimension. In particular, this leads us to the following, which summarizes Section˜3.
@@
Suppose $t\colon T\to S$ is a morphism of Noetherian schemes. Let $Y_{1}$ and $Y_{2}$ be proper and flat $S$-schemes. Denote by $\pi\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism. Consider $K\in D^{-}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
  1. (1)
    If $t$ is affine and $\Phi_{K}$ restricts to $D^{b}_{\operatorname{coh}}$ (resp. $\operatorname{Perf}$), then so does $\Phi_{\mathbf{L}\pi^{\ast}K}$ (see Lemma˜3.3). Additionally, the converse holds if $t$ is faithfully flat (see Corollary˜3.6).
  2. (2)
    Assume $t$ is affine and locally of finite tor-dimension.
    1. (a)
      If $\Phi_{K}$ restricts to $\operatorname{Perf}$, then $\mathbf{L}\pi^{\ast}$ preserves the adjunction of $\Phi_{K}$ with its right adjoint (see Proposition˜3.8).
    2. (b)
      If $\Phi_{K}$ restricts to $D^{b}_{\operatorname{coh}}$, then $\mathbf{L}\pi^{\ast}$ preserves the adjunction of $\Phi_{K}$ with its left adjoint (see Proposition˜3.12).
In short, Proposition˜1.1 describes the behavior of integral transforms under base change when restricted to the small categories, specifically addressing whether the preservation of adjoints occurs. Key instances where it applies are morphisms $t$ arising from field extensions or finite surjective morphisms to a Dedekind scheme (i.e. smooth projective curves).
Fully faithfulness & equivalences
With this understanding of adjointness under base change in hand, we turn to questions of fully faithfulness and equivalence. Particularly, regarding the preservation of adjointness and the restriction to smaller categories. Next, we shift gears to studying fully faithfulness and equivalences under base change; specifically, the ascent and descent of these properties. Before doing so, we review the relevant literature a bit more.
As mentioned earlier, the work of [HLS07, HLS09] studied these questions for fibrations whose base is of finite type over an algebraically closed field. A key ingredient in loc. cit. was to look at the behavior on special fibers. More recently, such questions have also been studied for smooth proper morphisms over a Noetherian base scheme (see [Kur24, Proposition/Definition 2.3]). However, smoothness assumptions restrict the range of examples one can consider, e.g. fibrations admitting singular fibers.
A result of Orlov addresses similar questions in the case of smooth varieties over a field. Specifically, for smooth projective varieties over a field, an integral transform is fully faithful or an equivalence if, and only if, the induced integral transform satisfies the same property after any base field extension [Orl02, Lemma 2.12]. The strategy of loc. cit. leverages ‘canonical morphisms’ between the kernels of integral transforms that describe the (co)units of adjunction. However, it remains unclear whether such canonical morphisms exist in the singular setting. This uncertainty provides further motivation for our work: we bypass the absence of such tools in the singular setting by developing different techniques.
This brings attention to the following, which summarizes Section˜4.
Theorem 1.2 (see Corollary˜4.7).
Consider affine morphism $t\colon T\to S$ locally of finite tor-dimension of Noetherian schemes whose target is semiseparated. Suppose $Y_{1}$ and $Y_{2}$ are proper and flat $S$-schemes. Denote by $\pi\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}\pi^{\ast}K}$.
In practice, many schemes we may consider are separated, and hence, semiseparated. As a pleasant consequence, this allows us to partially generalize Orlov’s result to the singular setting.
Corollary 1.3 (cf. [Orl02, Lemma 2.12]).
Consider proper schemes $Y_{1}$ and $Y_{2}$ over a field $k$. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{K_{L}}$ for any field extension $L/k$.
Our proofs require careful bookkeeping of how the (co)units of adjunctions for integral transforms behave under base change. However, additional subtleties also arise in this context. In particular, we need to work in the larger category $D_{\operatorname{qc}}$ in order to have a firm grasp on ascending fully faithfulness; see e.g. Proposition˜4.4 and Theorem˜4.6. Notably, the techniques we use are independent of those appearing in [Orl02, Lemma 2.12].
1.3. Why it matters
To conclude our work, we discuss the ramifications of the results above. Our results extend the literature on schemes of finite type over algebraically closed fields to the more general setting of arbitrary fields. They reveal a new interplay between the intuition of [HLS07, HLS09] concerning special fibers and faithfully flat descent along affine morphisms. In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes, broadening the scope to include arithmetically flavored situations of potential interest for future research.
@@ 1.3. Why it matters
For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, Fourier–Mukai partnership implies they are $k$-forms (see Proposition˜5.1). Here, a $k$-form means there exists a finite extension $L/k$ such that their base changes along $L/k$ become isomorphic. The result holds for singular curves and enriches algebraically closed case of [Spe23]. For fibrations of relative dimension one over a regular Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜5.2).
@@
Smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Proposition˜5.3). This subtlety arises especially over imperfect fields. Since regularity is characterized by $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$, one might expect this condition to persist after a change of field extensions, but there exist regular projective curves that become singular after such a base change (see e.g. [Kol11, Remark 16]).
@@
In the setting of projective fibrations, smoothness of the total space is also a derived invariant if the base is regular (see Corollary˜5.4). Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for regular Noetherian base schemes (see Proposition˜5.5).
In a related but distinct direction, we study a variation of Proposition˜5.5 in the case of derived equivalences. Particularly, we prove a version for the Gorenstein condition for arbitrary derived equivalences where the required functor need not be an integral transform. See Proposition˜5.6.
@@
Let $X$ be a Noetherian scheme. We work with the following triangulated categories:
@@
    $D_{\operatorname{coh}}^{b}(X)$ (resp. $D_{\operatorname{qc}}^{b}(X)$) is the (strictly full) subcategory of $D(X)$ consisting of complexes having bounded and coherent (resp. quasi-coherent) cohomology.
@@
At times if $X$ is affine, we abuse notation and write $D_{\operatorname{qc}}(R):=D_{\operatorname{qc}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories mentioned here and later.

### v3

@@ Abstract.
We study the behavior of integral transforms under base change. In particular, we establish a yoga of local algebra and fibers to test for derived equivalences or fully faithfulness via integral transforms. This generalizes a result of Orlov to singular varieties and strengthens several results in the literature by allowing arbitrary base fields. Additionally, it provides new insight into fibrations and their singularities in arithmetic settings (e.g. projective and flat schemes over a DVR).
@@
Fourier–Mukai partners, integral transforms, singular varieties, derived categories, fully faithfulness
@@ 1.1. What is known
The context for our work begins with Mukai [MUK81]. In that work, integral transforms were introduced to study triangulated equivalences between derived categories of abelian varieties. Since then, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g. [BRI02, KAW02]).
Much of this progress concerns smooth varieties, where geometric results can be obtained using categorical methods. However, extending these methods to singular varieties remains challenging, largely because bounded coherent complexes need not be perfect. Thus, an important task is to develop techniques for studying integral transforms in singular settings.
Recent work has begun to address these difficulties. Criteria determining when an integral transform is fully faithful or an equivalence have been established for varieties over algebraically closed fields, generalizing results from the smooth case [HLS07, HLS09] and building upon foundational work of Bondal–Orlov [A. 95]. Likewise, Bondal–Orlov’s reconstruction theorem for smooth projective varieties has been extended to projective curves [SPE23] and (anti-)Fano Gorenstein varieties [BAL11].
There is also growing interest in studying integral transforms in the relative setting. Specifically, let $Y_{1}$ and $Y_{2}$ be proper $S$-schemes, where $S$ is Noetherian. An integral transform is an exact functor
$\Phi_{K}:=\mathbf{R}(p_{2})_{\ast}(\mathbf{L}p_{1}^{\ast}(-)\otimes^{\mathbf{L}}K)\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$
where $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$ and $p_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ are the projection morphisms. If $\Phi_{K}$ is an equivalence, we say $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners (over $S$).
Such transforms naturally arise in the study of fibrations; that is, in situations involving proper and flat morphisms. See e.g. [HLS07, HLS09], which dealt with families whose special fibers have mild singularities. In particular, loc. cit. studied when an integral transform between locally projective fibrations $Y_{i}\to S$ is an equivalence for $S$ of finite type over an algebraically closed field. It was shown that the induced integral transform on special fibers (i.e. closed points) over $S$ controls being an equivalence or fully faithful.
@@ 1.2. What we do
In this work, we deepen this perspective by systematically exploiting special fibers. Specifically, we study the behavior of integral transforms under base change using techniques from local algebra.
@@
Neeman has developed a new framework called approximable triangulated categories (see e.g. [CNS25, NEE23]), which has resolved open conjectures in algebraic geometry [NEE21a, NEE24]. As a consequence, any integral transform $\Phi_{K}$ inducing an equivalence on $D_{\operatorname{qc}}$ restricts to equivalences on $D^{b}_{\operatorname{coh}}$ and $\operatorname{Perf}$, see Lemma˜2.10.
@@
Consequently, questions about equivalences may be reduced to the behavior of integral transforms on these smaller subcategories. Whether an integral transform restricts to these smaller categories is now understood [RIZ17, BAL09, DLM25]. These results shed light on the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
@@
However, the behavior of adjointness under base change is not yet fully understood. We focus is on base changes induced by certain morphisms. This leads us to the following.
Theorem 1.1.
Let $S$ be a Noetherian scheme. Suppose $Y_{1}$ and $Y_{2}$ are proper and flat $S$-schemes. Consider $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Then the following are equivalent:
  1. (1)
    $K$ is relatively perfect over $Y_{1}$ (resp. $Y_{2}$)
  2. (2)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for any affine morphism $t$ from a Noetherian scheme
  3. (3)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(k)$ (resp. $Y_{2}\times_{S}\operatorname{Spec}(k)$) for all morphisms $t\colon\operatorname{Spec}(k)\to S$ from a field
  4. (4)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}\operatorname{Spec}(\kappa(p))$ (resp. $Y_{2}\times_{S}\operatorname{Spec}(\kappa(p))$) for every closed point $p\in S$ with associated closed immersion $t\colon\operatorname{Spec}(\kappa(p))\to S$
  5. (5)
    $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{1}\times_{S}T$ (resp. $Y_{2}\times_{S}T$) for some affine surjective morphism $t\colon T\to S$.
Here, $t^{\prime}\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ is the natural morphism. In fact, in either case the corresponding left or right adjoints are preserved.
This follows from Theorem˜3.6 together with Propositions˜3.8 and 3.11. In short, Theorem˜1.1 describes how integral transforms behave under base change when restricted to $\operatorname{Perf}$ or $D^{b}_{\operatorname{coh}}$. Importantly, it addresses the preservation of adjoints under base change. To the best of our knowledge, this result is new. In particular, it applies in the presence of singularities on the $Y_{i}$. An important instance where the result is applicable is $t$ arising from field extensions.
Fully faithfulness & equivalences
Next, we shift gears to studying fully faithfulness and equivalences under base change. Specifically, the ascent and descent of these properties. Before doing so, we briefly review the relevant literature.
As mentioned earlier, [HLS07, HLS09] studied these questions for fibrations whose base is of finite type over an algebraically closed field. A key ingredient in their work was to look at the behavior on special fibers. More recently, such questions have also been studied for smooth proper morphisms over a Noetherian base scheme (see [KUR24, Proposition/Definition 2.3]). However, smoothness assumptions restrict the range of examples one can consider, e.g. fibrations admitting singular fibers.
A result of Orlov addresses similar questions in the case of smooth varieties over a field. Specifically, for smooth projective varieties over a field, an integral transform is fully faithful or an equivalence if, and only if, the induced integral transform satisfies the same property after any base field extension [ORL02, Lemma 2.12]. The strategy of loc. cit. ises ‘canonical morphisms’ between the kernels of integral transforms that describe the (co)units of adjunction. However, it remains unclear whether such canonical morphisms exist in the singular setting. This uncertainty provides further motivation for our work: we bypass the absence of such tools in the singular setting by developing different techniques.
This brings attention to the following.
Theorem 1.2.
Let $S$ be a Noetherian scheme. Suppose $Y_{1}$ and $Y_{2}$ are proper and flat $S$-schemes. Consider $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ which is relatively perfect over each $Y_{i}$. Then the following are equivalent:
  1. (1)
    $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$
  2. (2)
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for any affine morphism $t$
  3. (3)
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for every morphism $t\colon\operatorname{Spec}(k)\to S$ from a field
  4. (4)
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for every closed point $p\in S$ with associated closed immersion $t\colon\operatorname{Spec}(\kappa(p))\to S$
  5. (5)
    $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for some affine surjection $t$.
Here, $t^{\prime}\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ is the natural morphism induced by $t\colon T\to S$.
This appears later as Theorem˜4.13. Notably, it generalizes [HLS09, Proposition 2.15], which only proved the statement for closed points. As a consequence, this allows us to generalize Orlov’s result to the singular setting (cf. [ORL02, Lemma 2.12]).
Corollary 1.3.
Consider proper schemes $Y_{1}$ and $Y_{2}$ over a field $k$. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{k}Y_{2})$ be relatively perfect over each $Y_{i}$. Then $\Phi_{K_{L}}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$ for any field extension $L/k$ if, and only if, $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$.
Our proofs require careful bookkeeping of how the (co)units of adjunction for integral transforms behave under base change. See Section˜4 for details. Notably, the techniques we use are independent of those appearing in [ORL02, Lemma 2.12].
1.3. What is next
We conclude by discussing several consequences of the results above. Our results extend the literature on schemes of finite type over algebraically closed fields to the more general setting of arbitrary fields. They reveal a new interplay between the intuition of [HLS07, HLS09] concerning special fibers and faithfully flat descent along affine morphisms. In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes. This broadens the scope to include arithmetically flavored situations of potential interest for future research.
For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, we show Fourier–Mukai partnership implies they are $k$-forms. Particularly:
@@ 1.3. What is next
Here, a $k$-form means there exists a finite extension $L/k$ such that their base changes along $L/k$ become isomorphic. The result holds for singular curves and extends the algebraically closed case of [SPE23]. For fibrations of relative dimension one over a Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜5.1). It would be interesting to find a prime or mixed characteristic analog of these results.
@@
Next, we show that smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Lemma˜5.2). This subtlety arises especially over imperfect fields. Since regularity is characterized by $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$, one might expect this condition to persist after a change of field extensions, but there exist regular projective curves that become singular after such a base change (see e.g. [KOL11, Remark 16]).
In the setting of fibrations, smoothness of the total space is also a derived invariant:
Proposition 1.5.
Let $f_{i}\colon Y_{i}\to S$ be proper flat morphisms to a Noetherian scheme. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is smooth if, and only if, $f_{2}$ is smooth.
@@
Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for regular Noetherian base schemes. Specifically:
Proposition 1.6.
Let $f_{i}\colon Y_{i}\to S$ be projective flat morphisms with geometrically integral fibers to a Noetherian scheme. Assume that $Y_{1}$ and $Y_{2}$ are Fourier–Mukai partners over $S$. Then $f_{1}$ is Cohen–Macaulay (resp. Gorenstein) if, and only if, $f_{2}$ satisfies the same condition.
Under additional assumptions, Proposition˜1.6 can be strengthened in the Gorenstein case. See Proposition˜5.3 for details. Its proof relies on techniques that differ from those in Proposition˜1.6.
Finally, we comment on potential generalizations of our work. The authors are currently investigating stack theoretic consequences of these results. Under mild assumptions, the work of [AJS23] extends to suitable algebraic stacks. At present, it appears that much of our work carries over to this setting with a bit of effort. We also plan to explore deeper applications to Fourier–Mukai theory for moduli spaces using the tools developed here.
@@
Let $X$ be a scheme. We work with the following triangulated categories:
@@
    If $X$ is Noetherian, then $D_{\operatorname{coh}}^{b}(X)$ (resp. $D_{\operatorname{qc}}^{b}(X)$) is the (strictly full) subcategory of $D(X)$ consisting of complexes having bounded and coherent (resp. quasi-coherent) cohomology.
@@
At times if $X$ is affine, we might abuse notation and write $D^{b}_{\operatorname{coh}}(R):=D^{b}_{\operatorname{coh}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories mentioned here and later.